OpenStudy (anonymous):

find any intercepts. y=x^2+3x / (3x+1)^2

OpenStudy (anonymous):

please!

OpenStudy (anonymous):

\[y = \frac{x^2 + 3x}{(3x + 1)^2}\]

OpenStudy (anonymous):

yes, I thought you start off setting y to 0, then finding those intercepts?

OpenStudy (anonymous):

but I'm not sure how to simplify the right

OpenStudy (hba):

ok when x =0 u fin b and when y=o u find a

OpenStudy (anonymous):

Just factor the top: \[y = \frac{x(x + 3)}{(3x + 1)^2}\]

OpenStudy (anonymous):

You can find the x-int when x = 0, so that equation will become 0 when the numerator is 0. So find the solutions in the numerator, and you'll have your x-intercepts. Remember the domain also because the denominator can't be 0 or it'll become undefined

OpenStudy (anonymous):

but how would I find the y?

OpenStudy (anonymous):

Just plug in 0 for the x-values and the number you get is the y-intercept

OpenStudy (anonymous):

So plug in 0 for all x values.. then what I get is the y? let me try and tell you what i get

OpenStudy (anonymous):

is it 0?

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

So y is (0,0)

OpenStudy (anonymous):

how would i find x?

OpenStudy (anonymous):

Set the numerator I gave you above to 0 and find each x-value that will satisfy the value to be 0

OpenStudy (anonymous):

can u explain?

OpenStudy (anonymous):

i thought you set y to 0..then solved

OpenStudy (anonymous):

The numerator factored is x(x + 3), right? So set each x to 0 and you have two equations to solve: x = 0 and x + 3 = 0 Those will be your x-intercepts

OpenStudy (anonymous):

Ok what about the denominator

OpenStudy (anonymous):

so I have y= (0,0) and x = 0, -3

OpenStudy (anonymous):

The denominator doesn't matter because whatever the denominator is, it won't make the equation 0. Just make sure it doesn't become undefined, that's all

OpenStudy (anonymous):

Ok so is what I have above correct? like is that the answer

OpenStudy (anonymous):

Those are your 3 intercepts, yes